OpenStudy (anonymous):

The base of an isosceles triangle is half the length of each of its congruent sides. If the perimeter of the triangle is 120 inches, what is the length of the base? A. 24 inches B. 16 inches C. 6 inches D. 12 inches

OpenStudy (anonymous):

I'll give medal and fan!!!! PLEAS!

OpenStudy (anonymous):

Please help\

OpenStudy (michele_laino):

Hint: if I call with L the length of each congruent sides, then the length of the base is L/2, so you have to solve this equation: \[L+L+\frac{ L }{ 2 }=120\] please what is L?

OpenStudy (anonymous):

hold on i'm thinking more hint??

OpenStudy (anonymous):

please help

OpenStudy (michele_laino):

here is another hint: \[L+L+\frac{ L }{ 2 }=\frac{ 5 }{ 2 }L\] so your equation, becomes: \[\frac{ 5 }{ 2 }L=120\]

OpenStudy (michele_laino):

so, what is L?

OpenStudy (michele_laino):

please, keep in mind that you have to find the value of L/2

OpenStudy (anonymous):

I'll do my best

OpenStudy (anonymous):

I sort of understand so i'll give you your medal and I already fanned you so..... yeah bye!!! Oh and THANKS A BUNCH!!

OpenStudy (anonymous):

Medal given

OpenStudy (michele_laino):

thanks! :)

OpenStudy (michele_laino):

so, please do you know how to solve the subsequent equation? \[\frac{ 5 }{ 2 }L = 120\]